OpenStudy (anonymous):

find the formula for df^-1/dx. f(x)=(7-x)^3

OpenStudy (anonymous):

what's d?

OpenStudy (ujjwal):

can you find \(f^{-1}(x)\) ?

OpenStudy (anonymous):

oh. ok f(x) = (7-x)^3 y = (7-x)^3 switch y and x x = (7 - y)^3 third root of x = 7 - y -y = 3rootx - 7 so \[f \prime(x) = \sqrt[3]{x} - 7\]

OpenStudy (anonymous):

Thanks for the help but that's not an option on the homework for an answer.

OpenStudy (anonymous):

The answer should be y = - 3 root x +7

OpenStudy (anonymous):

It is the same thing as the above poster's answer except the person forgot to divide by -1

OpenStudy (anonymous):

oh yeah, lol sorry about that :P

OpenStudy (anonymous):

\[7-x ^{1/3}\]\[-3(7-x)^{2}\]\[x ^{2/3}\]\[-1/3x ^{2/3}\] are the answer options on the homework.

OpenStudy (anonymous):

it's 7 - x^1/3

OpenStudy (anonymous):